Having trouble. Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.
4x – 5y = -7
-8x + 10y = 14

Answer by Alan3354(69443)

4x – 5y = -7
-8x + 10y = 14
------------------
The 2nd eqn is -2 times the first
--> dependent, infinite # of solutions, no unique solution.